# Peeking with PEAK: Sequential, Nonparametric Composite Hypothesis Tests for Means of Multiple Data Streams

### v2

@@ Abstract
We propose a novel nonparametric sequential test for composite hypotheses for means of multiple data streams. Our proposed method, peeking with expectation-based averaged capital (PEAK), builds upon the testing-as-betting framework and provides a non-asymptotic $\alpha$-level test across any stopping time. PEAK is computationally tractable and efficiently rejects hypotheses that are incorrect across all potential distributions that satisfy our nonparametric assumption, enabling joint composite hypothesis testing on multiple streams of data. We numerically validate our theoretical findings under the best arm identification and threshold identification in the bandit setting, illustrating the computational efficiency of our method against state-of-the-art testing methods.
@@ 1 Introduction
Sequential experiments have important applications in various fields to optimize decision-making, including better content recommendation (Bouneffouf et al.,, 2012), personalized learning (Clement et al.,, 2015; Cai et al.,, 2021), and optimized digital interventions (Figueroa et al.,, 2021; Forman et al.,, 2019; Liao et al.,, 2020; Piette et al.,, 2022; Trella et al.,, 2022, 2023; Yom-Tov et al.,, 2017). Providing informative and anytime-valid feedback during sequential experiments could potentially lead cost reduction through early stopping (Liang and Bojinov,, 2023) and improved experiment outcomes. For example, in digital interventions, prompt “expert” feedback has been shown to increase user engagement, resulting in better treatment outcomes (Sharpe et al.,, 2017; Yardley et al.,, 2016). In this work, we aim to provide anytime-valid inference in sequential experiments for a task of interest (e.g., identifying the best arm/treatment in the bandit/reinforcement learning setting), where data collection adapts based on previously observed outcomes. This problem can be formulated as a sequential hypothesis testing problem (Ramdas et al.,, 2022). The problem necessitates special inferential tools, since naively repeatedly applying conventional inference such as $t$-tests at every time point will always at some point detect an effect even when none exists.
In the settings where sample sizes are limited and multiple arms are to be compared, the ideal testing tools for sequential and possibly adaptive experiments should fulfill the following desiderata:
@@
Various sequential testing methods focus on asymptotic guarantees (Bibaut et al.,, 2022; Robbins and Siegmund,, 1970; Waudby-Smith et al.,, 2021; Woong Ham et al.,, 2023), where approximately valid (and powered) inference only start after sufficiently large sample sizes. While these methods appeal to other desiderata such as asymptotic calibration of type-I errors and optimal power, they may not be appropriate when inference is desired for small sample sizes, reflected in our desideratum (a).
@@
Recent works utilizing $e$-processes for non-asymptotic anytime-valid inference fall into two groups. The first involves Chernoff-based approach (Howard et al., 2021a, ), which relies on sub-$\psi$ assumptions (such as sub-gaussian/sub-bernoulli) combined with peeling (Cappé et al.,, 2013) or mixture techniques (de la Peña et al.,, 2004) . The second involves the testing-as-betting framework (Waudby-Smith and Ramdas,, 2023). In the bounded observation setting, both empirical (Waudby-Smith and Ramdas,, 2023) and theoretical (Shekhar and Ramdas,, 2023) evidence indicate that the latter outperforms the former. However, these approaches suffer from computational inefficiency – computing confidence sequences and conducting composite tests require performing a grid search over the hypothesis space of means, which grows in dimension as the number of data streams increases. Thus, while the current testing schemes proposed in Waudby-Smith and Ramdas, (2023) satisfy both desiderata (a), (b), they fall short of satisfying (c).
Contributions.  In this paper, we provide a novel nonparametric sequential test for composite hypotheses on bounded means that aim to achieve desiderata (a)-(c). We build upon the testing-as-betting framework, and establish theoretical guarantees for our procedure. Our testing scheme significantly reduces the computational overhead compare to other testing-as-betting sequential tests, enabling tractable testing for popular hypotheses (e.g., best-arm identification, threshold identification) in the multi-armed bandit setting. Our method shares the computational convenience of tail-inequality-based sequential tests, while maintaining the performance of the state-of-the-art testing-as-betting methods.
@@
Outline.  In Section 2, we define $e$-values and introduce the concept of $e$-power, which serves as the natural analogue of power in settings involving optional stopping and continuation. Next, we present the testing-as-betting framework by Waudby-Smith and Ramdas, (2023). In Section 3, we introduce our testing procedure in the single arm/stream setting, and provide theoretical results demonstrating that this test (i) controls type I error, (ii) is of power 1, and (iii) forms convex/continuous confidence intervals. In addition, we quantify the asymptotic suboptimality of our test. In Section 4, we provide a new test based on averaging evidence across arms, and show that this testing approach maintains computational tractability while strictly dominating a union-bound approach. In Section 5, we numerically test our sequential testing scheme. In the single-arm setting, we demonstrate that the resulting confidence sequences are comparable to those produced by the state-of-the-art schemes. We additionally evaluate our testing scheme in both the best-arm identification and threshold identification settings, illustrating that our scheme yields improved expected stopping time when compared with existing approaches.
@@ 2 Problem Formulation and Related Work
We assume the nonparametric setting, making no further assumptions about the data-generating process (DGP), allowing for continuous and discrete distributions over $[0,1]$. For each hypothesis $m_{a}\in[0,1]$ of arm $a$, let $\mathcal{P}(m_{a})$ denote the set of all distributions on $[0,1]$ with mean $m_{a}$. Each hypothesis vector $m=(m_{1},...,m_{W})\in[0,1]^{W}$ denotes the composite hypothesis $\cap_{a\in[W]}\{P_{a}\in\mathcal{P}(m_{a})\}$, in which we fix the distribution means at $m$ but otherwise allow for arbitrary distributions over $[0,1]$ for each arm. We use $R\subseteq[0,1]^{W}$ to denote the set of composite hypotheses of means (corresponding to regions of the hypothesis space) that we are interested in testing at time $t\in\mathbb{N}$. We defer further discussion on $R$ to Section 2.2.
A sequential test $T_{t}(R)$ for a composite hypothesis $R$ is a binary-valued $\mathcal{F}_{t}$-measurable variable, with value 1 indicating rejection of $R$. Our goal is to develop a computationally tractable sequential test $T_{t}(R)$ that maintains a $\alpha$-level type I error rate, i.e., whenever $P_{a}\in\mathcal{P}(m_{a})$ for all $a$ for some $m\in R$, we have
@@
The test we developed will follow from the $e$-value and the testing-as-betting frameworks (Waudby-Smith and Ramdas,, 2023). We formally introduce these concepts below.
2.1 $e$-Processes and Testing-as-Betting Framework
@@
If $K_{t}$ is an $e$-process for hypothesis $\mathcal{H}$, the $\alpha$-level sequential test is to simply reject $\mathcal{H}$ if $K_{t}$ ever exceeds the threshold $1/\alpha$. The design of $e$-value-based tests directly leads to the interpretation of $K_{t}$ as evidence against the null: we reject $\mathcal{H}$ when $K_{t}$ is sufficiently large. The analogue of power of an $e$-process based test, called the growth rate or $e$-power (Vovk and Wang,, 2024), corresponds to this intuition.
@@
For an $e$-value $E$ with respect to null hypothesis $P$, the growth rate of $E$ with respect to an alternative hypothesis $Q$ is given by $\mathbb{E}_{Q}[\log(E)]$.
In the case of $e$-process $K_{t}$, the $e$-power is simply the sum of the expected log-values for each conditional $e$-variable $E_{t}$. Maximizing the $e$-power intuitively means maximizing the evidence against the null when the null is misspecified; under certain conditions, this directly corresponds to minimizing the expected stopping time in the i.i.d. single arm setting (Grünwald et al.,, 2023; ter Schure et al.,, 2023). Moreover, for simple null hypothesis $P$ and simple alternative hypothesis $Q$ where $Q$ is absolutely continuous with respect to $P$, the likelihood ratio $dQ/dP$ maximizes the $e$-power (Lemma 2.1 of Vovk and Wang,, 2024), aligning directly with the same notion of power in classical parametric sequential tests (Wald and Wolfowitz,, 1948).
Testing-as-Betting Framework.   In the single arm setting for bounded means, with null hypothesis $R=m$, the test we develop can be seen as a special case of the more general testing-as-betting framework (Waudby-Smith and Ramdas,, 2023), which constructs tests based on sizing bets. Specifically, the testing-as-betting framework constructs the $e$-process $M_{t}(m)$ as follows:
@@
for some $\lambda_{t}(m)\in[-1/(1-m),1/m]$, where $\lambda_{t}(m)$ is $\mathcal{F}_{t-1}$-measurable. [^1] Let $1/m=\infty$ when $m=0$, and $-\frac{1}{1-m}=-\infty$ when $m=1$. The sequence $\lambda_{t}(m)$ can be interpreted as bets against the belief of whether $\mu$ is above $m$, with the sign indicting direction of the belief and the magnitude indicting the strength of the belief (i.e., how much of our current evidence $M_{t}(m)$ we are willing to put at stake). We discuss our novel choice of $\lambda_{t}(m)$ in Section 3, using the single-arm case as a guiding example.
Figure 1: Common composite hypotheses of interest for pure exploration bandit problems with two arms ($A=2$). From left to right: thresholding bandit problem, best-arm-identification, best arm above threshold.
@@ 2.2 Hypothesis Testing of Means for Multiple Streams
Other potential hypotheses of interest include determining if each stream’s mean is above/below a certain threshold (Kano et al.,, 2018). For such objectives, the hypotheses form convex regions of the hypothesis space $[0,1]^{W}$, shown in Figure 1. In Section 4.2, we show our test enables computationally tractable testing for such convex regions.
Another work that directly considers joint sequential tests across arms is Kaufmann and Koolen, (2021), which uses a mixture-based approach to construct $e$-processes for a single arm, and combines this evidence using the product of evidence across arms. While their objective aligns with ours, the setting of their work differ drastically: the parameters of the data streams are assumed to belong to 1-dimensional exponential families, rather than our nonparametric setting.
The proposed testing-as-betting methods in Waudby-Smith and Ramdas, (2023) suffer from computational inefficiency when generalized to the a multi-armed case. In the single stream setting, these tests require a search over the hypothesis space $[0,1]$, using a grid search to rule out hypotheses $m$ sequentially. In the multiple stream setting, the size of the grid search grows exponentially with the dimension $W$. In contrast, our tests remove the need for grid search.
@@ 3 Tests for a Single Stream of Data
The capital process $K_{t}(m)$, and the corresponding log capital process $l_{t}(m)$ are defined as as follows:
@@
Let $e_{i}(m)=1+\frac{(\hat{\mu}_{i-1}-m)(X_{i}-m)}{c}$. Then, (1) implies that $e_{i}(m)$ is an $e$-variable. The captial process $K_{t}(m)$ can be written as $K_{t}(m)=\prod_{i=1}^{t}e_{i}(m)$. By Definition 1, the capital process $K_{t}(m)$ forms an $e$-process for hypothesis $m$. We note that $K_{t}(m)$, (6), is a special case of the nonnegative martingale construction $M_{t}(m)$, (2), with $\mathcal{F}_{t-1}$-measurable betting sequence $\{\lambda_{t}(m)=\frac{(\hat{\mu}_{t-1}-m)}{c}\}_{t\in\mathbb{N}}$. This betting scheme is intuitive: we place our bet according to the sign of $\hat{\mu}_{t-1}-m$ (a plug-in estimate of the sign of $X_{t}-m_{a}$) and base the bet’s magnitude on the absolute distance between $\hat{\mu}_{t-1}$, our running estimate of $X_{t}$, and hypothesis $m$.
@@
$T_{t}(m,\alpha)=\mathbf{1}[\max_{1\leq i\leq t}K_{i}(m)\geq 1/\alpha].$ (6)
@@
For all $c\geq 1/4$, $T_{t}(m,\alpha)$ defines a sequential test with type I error $\alpha$ for null hypothesis $m$, i.e., if $\mu=m$ (equivalently, $P_{1}\in\mathcal{P}(m)$), $\mathbb{P}(\exists t\in\mathbb{N}:T_{t}(m)=1)\leq\alpha$.
@@
The convexity of our confidence sequence $C_{t}(\alpha)$ follows directly from Proposition 1 (proof in Appendix A.6): at time $t=0$, $C_{0}(\alpha)=[0,1]$; for $t\in\mathbb{N}$, each test results in a convex set; repeated intersections $C_{t}(\alpha)$ is also convex. Note that this property does not generally hold in testing-as-betting schemes by Waudby-Smith and Ramdas, (2023).
While Theorems 1 and 2 guarantee that $\mu$ will be identified in our test, they do not provide additional insights. To better understand the behavior of our test, in Section 3.1 Lemma 1, we characterize the asymptotic $e$-power/growth rate (Definition 2) of our process, $K_{t}(m)$, which defines the expected linear rate of increase of $l_{t}(m)$ as $t\rightarrow\infty$. We discuss the selection of parameter $c$ in Lemma 2. In Lemma 3, we establish that the worst-case instance of $e$-power is strictly positive for $m\neq\mu$. Finally, given that our test is non-parametric, we compare the asymptotic growth rate of our test in the parametric setting to the best-achievable growth rate in Lemma 4, establishing that the $e$-power of our test is near-optimal in the parametric setting.

### v3

@@ Abstract
We propose a novel nonparametric sequential test for composite hypotheses for means of multiple data streams. Our proposed method, peeking with expectation-based averaged capital (PEAK), builds upon the testing-by-betting framework and provides a non-asymptotic $\alpha$-level test across any stopping time. Our contributions are two-fold: (1) we propose a novel betting scheme and provide theoretical guarantees on type-I error control, power, and asymptotic growth rate/$e$-power in the setting of a single data stream; (2) we introduce PEAK, a generalization of this betting scheme to multiple streams, that (i) avoids using wasteful union bounds via averaging, (ii) is a test of power one under mild regularity conditions on the sampling scheme of the streams, and (iii) reduces computational overhead when applying the testing-as-betting approaches for pure-exploration bandit problems. We illustrate the practical benefits of PEAK using both synthetic and real-world HeartSteps datasets. Our experiments show that PEAK provides up to an 85% reduction in the number of samples before stopping compared to existing stopping rules for pure-exploration bandit problems, and matches the performance of state-of-the-art sequential tests while improving upon computational complexity.
@@ 1 Introduction
Sequential experiments have important applications in various fields to optimize decision-making, including better content recommendation (Bouneffouf et al.,, 2012), personalized learning (Clement et al.,, 2015; Cai et al.,, 2021), and optimized digital interventions (Figueroa et al.,, 2021; Forman et al.,, 2019; Liao et al.,, 2020; Piette et al.,, 2022; Trella et al.,, 2022, 2023; Yom-Tov et al.,, 2017). Providing informative and anytime-valid feedback during sequential experiments could potentially lead to cost reduction through early stopping (Liang and Bojinov,, 2023) and improved experiment outcomes. For example, in digital interventions, prompt “expert” feedback has been shown to increase user engagement, resulting in better treatment outcomes (Sharpe et al.,, 2017; Yardley et al.,, 2016). In this work, we aim to provide anytime-valid inference in sequential experiments for a task of interest (e.g., identifying the best arm/treatment in the bandit/reinforcement learning setting), where data collection adapts based on previously observed outcomes. The problem necessitates special inferential tools since naively repeatedly applying conventional inference such as $t$-tests at every time point will always at some point detect an effect even when none exists.
In settings where sample sizes are limited and multiple arms are to be compared, the ideal testing tools for sequential and possibly adaptive experiments should fulfill the following desiderata:
@@
Various sequential testing methods focus on asymptotic guarantees (Bibaut et al.,, 2022; Robbins and Siegmund,, 1970; Waudby-Smith et al.,, 2021; Woong Ham et al.,, 2023), where approximately valid (and powered) inference only starts after sufficiently large sample sizes. While these methods appeal to other desiderata such as asymptotic calibration of type-I errors and optimal power, they may not be appropriate when inference is desired for small sample sizes, reflected in our desideratum (a).
@@
Recent works utilizing $e$-processes for non-asymptotic anytime-valid inference fall into two groups. The first involves Chernoff-based approach (Howard et al., 2021a, ), which relies on sub-$\psi$ assumptions (such as sub-gaussian/sub-bernoulli) combined with peeling (Cappé et al.,, 2013) or mixture techniques (de la Peña et al.,, 2004). The second involves the testing-by-betting framework (Waudby-Smith and Ramdas,, 2023). In the bounded observation setting, both empirical (Waudby-Smith and Ramdas,, 2023) and theoretical (Shekhar and Ramdas,, 2023) evidence indicate that the latter outperforms the former. However, these approaches suffer from computational inefficiency – computing confidence sequences and conducting composite tests require performing a grid search over the hypothesis space of means, which grows in dimension as the number of data streams increases. Thus, while the current testing schemes proposed in Waudby-Smith and Ramdas, (2023) satisfy both desiderata (a), (b), they fall short of satisfying (c).
Contributions.  In this paper, we provide a novel nonparametric sequential test for composite hypotheses on bounded means that aim to achieve desiderata (a)-(c). We build upon the testing-by-betting framework and establish theoretical guarantees for our procedure. Our contributions are two-fold: (1) we propose a novel betting scheme, and provide theoretical guarantees on type-I error control, power, and asymptotic growth rate/$e$-power that have not been shown for other adaptive testing-by-betting methods in the setting of a single data stream; (2) we introduce PEAK, a generalization of this betting scheme to multiple streams, that (i) avoids using wasteful union bounds via averaging, (ii) is a test of power one under mild regularity conditions on the sampling scheme of the streams, and (iii) reduces computational overhead when applying the testing-by-betting approaches for pure-exploration bandit problems.
@@
Outline.  In Section 2, we define $e$-values and introduce the concept of $e$-power, which serves as the natural analogue of power in settings involving optional stopping and continuation. Next, we present the testing-by-betting framework by Waudby-Smith and Ramdas, (2023). In Section 3, we introduce our testing procedure in the single arm/stream setting, and provide theoretical results demonstrating that this test (i) controls type-I error, (ii) is of power 1, and (iii) forms convex/continuous confidence intervals. In addition, we quantify the asymptotic suboptimality of our test. In Section 4, we provide a new test based on averaging evidence across arms, and show that this testing approach maintains computational tractability while strictly dominating a union-bound approach. In Section 5, we numerically test our sequential testing scheme for both single-arm and multi-arm settings on both simulated data and the mHealth HeartSteps dataset (Liao et al.,, 2020), demonstrating the efficacy of our approach.
@@ 2 Problem Formulation and Related Work
We assume the nonparametric setting, making no further assumptions about the data-generating process (DGP), allowing for continuous and discrete distributions over $[0,1]$. For each hypothesis $m_{a}\in[0,1]$ of arm $a$, let $\mathcal{P}(m_{a})$ denote the set of all distributions on $[0,1]$ with mean $m_{a}$. Each hypothesis vector $m=(m_{1},...,m_{W})\in[0,1]^{W}$ denotes the composite hypothesis $\cap_{a\in[W]}\{P_{a}\in\mathcal{P}(m_{a})\}$, in which we fix the distribution means at $m$ but otherwise allow for arbitrary distributions over $[0,1]$ for each arm. [^1] The minimal assumption necessary on the arm distributions for all results in this document can be found in Section 2, paragraph 2 of Waudby-Smith and Ramdas,, 2023. It only requires stationary of the expectations for each arm conditional on the previous history. We use $R\subseteq[0,1]^{W}$ to denote the set of composite hypotheses of means (corresponding to regions of the hypothesis space) that we are interested in testing at time $t\in\mathbb{N}$. We defer further discussion on $R$ to Section 2.2.
A sequential test $T_{t}(R)$ for a composite hypothesis $R$ is a binary-valued $\mathcal{F}_{t}$-measurable variable, with value 1 indicating rejection of $R$. Our goal is to develop a computationally tractable sequential test $T_{t}(R)$ that maintains a $\alpha$-level type-I error rate, i.e., whenever $P_{a}\in\mathcal{P}(m_{a})$ for all $a$ for some $m\in R$, we have
@@
The test we developed will follow from the $e$-value and the testing-by-betting frameworks (Waudby-Smith and Ramdas,, 2023). We formally introduce these concepts below.
2.1 $e$-Processes and Testing-by-Betting Framework
@@
If $K_{t}$ is an $e$-process for hypothesis $\mathcal{H}$, the $\alpha$-level sequential test is to reject $\mathcal{H}$ if $K_{t}$ ever exceeds the threshold $1/\alpha$. The design of $e$-value-based tests directly leads to the interpretation of $K_{t}$ as evidence against the null: we reject $\mathcal{H}$ when $K_{t}$ is sufficiently large. The analogue of power for an $e$-process-based test, called the growth rate or $e$-power (Vovk and Wang,, 2024), corresponds to this intuition.
@@
For an $e$-value $E$ concerning the null hypothesis $P$, the growth rate of $E$ with respect to an alternative hypothesis $Q$ is given by $\mathbb{E}_{Q}[\log(E)]$.
In the case of $e$-process $K_{t}$, the $e$-power is the sum of the expected log values for each conditional $e$-variable $E_{t}$. Maximizing the $e$-power intuitively means maximizing the evidence against the null when the null is misspecified; under certain conditions, this directly corresponds to minimizing the expected stopping time in the i.i.d. single arm setting (Grünwald et al.,, 2023; ter Schure et al.,, 2023). Moreover, for simple null hypothesis $P$ and simple alternative hypothesis $Q$ where $Q$ is absolutely continuous with respect to $P$, the likelihood ratio $dQ/dP$ maximizes the $e$-power (Lemma 2.1 of Vovk and Wang,, 2024), aligning directly with the same notion of power in classical parametric sequential tests (Wald and Wolfowitz,, 1948).
Testing-by-Betting Framework.   In the single-arm setting for bounded means, with null hypothesis $R=m$, the test we develop can be seen as a special case of the more general testing-by-betting framework (Waudby-Smith and Ramdas,, 2023), which constructs tests based on sizing bets. Specifically, the testing-by-betting framework constructs the $e$-process $M_{t}(m)$ as follows:
@@
for some $\lambda_{t}(m)\in[-1/(1-m),1/m]$, where $\lambda_{t}(m)$ is $\mathcal{F}_{t-1}$-measurable. [^2] Let $1/m=\infty$ when $m=0$, $-\frac{1}{1-m}=-\infty$ when $m=1$. The sequence $\lambda_{t}(m)$ can be interpreted as bets against the belief of whether $\mu$ is above $m$, with the sign indicating the direction of the belief and the magnitude indicating the strength of the belief (i.e., how much of our current evidence $M_{t}(m)$ we are willing to put at stake). We discuss our novel choice of $\lambda_{t}(m)$ in Section 3, using the single-arm case as a guiding example.
The proposed testing-by-betting methods (i.e., choice of $\lambda_{t}(m)$) in Waudby-Smith and Ramdas, (2023) suffer from computational inefficiency when generalized to the multi-armed case. In the single-stream setting, these tests require a search over the hypothesis space $[0,1]$, using grid search to rule out hypotheses $m$ sequentially. In the multi-stream setting, the size of the grid search grows exponentially with the dimension $W$. Our choice of sequential test, even with multiple streams, removes the need for grid search and allows various convex optimization approaches to be used.
@@ 2.2 Hypothesis Testing of Means for Multiple Streams
Other potential hypotheses of interest include determining if each stream’s mean is above/below a certain threshold (Kano et al.,, 2019). For such objectives, the hypotheses form convex regions of the hypothesis space $[0,1]^{W}$. In Section 4.2, we show our test enables computationally tractable testing for such convex regions.
Another work that directly considers joint sequential tests across arms is Kaufmann and Koolen, (2021), which uses a mixture-based approach to construct $e$-processes for a single arm and combines this evidence using the product of evidence across arms. While their objective aligns with ours, the setting of their work differ drastically: the parameters of the data streams are assumed to belong to 1-dimensional exponential families, rather than our nonparametric setting.
@@ 3 Tests for a Single Stream of Data
The capital process $K_{t}(m)$ and the corresponding log capital process $l_{t}(m)$ are defined as follows:
@@
Let $e_{i}(m)=1+\frac{(\hat{\mu}_{i-1}-m)(X_{i}-m)}{c}$. Then, (1) implies that $e_{i}(m)$ is an $e$-variable. The capital process $K_{t}(m)$ can be written as $K_{t}(m)=\prod_{i=1}^{t}e_{i}(m)$. By Definition 1, the capital process $K_{t}(m)$ forms an $e$-process for hypothesis $m$. We note that $K_{t}(m)$, (6), is a special case of the nonnegative martingale construction $M_{t}(m)$, (2), with $\mathcal{F}_{t-1}$-measurable betting sequence $\{\lambda_{t}(m)=\frac{(\hat{\mu}_{t-1}-m)}{c}\}_{t\in\mathbb{N}}$. This betting scheme is intuitive: we place our bet according to the sign of $\hat{\mu}_{t-1}-m$ (a plug-in estimate of the sign of $X_{t}-m_{a}$) and base the bet’s magnitude on the absolute distance between $\hat{\mu}_{t-1}$ (our running estimate of $X_{t}$) and hypothesis $m$.
We note that our choice of $\lambda_{t}(m)=\frac{\hat{\mu}_{t-1}-m}{c}$ differs from all existing betting schemes. The closest betting scheme to our choice of $\{\lambda_{t}(m)\}_{t\in\mathbb{N}}$ is the AGRAPA method (Appendix B, Waudby-Smith and Ramdas, 2023), which sets $\lambda_{i}(m)=\max(\frac{-l}{1-m},\min\{\frac{l}{m},\frac{\hat{\mu}_{i-1}-m}{\sigma_{i-1}^{2}+(\hat{\mu}_{i-1}-m)^{2}}\})$ for some $l\leq 1$ and a $\mathcal{F}_{i-1}$-measurable estimate of the variance, $\hat{\sigma}_{i-1}^{2}$. We discuss the differences between our $\lambda$ scheme and AGRAPA in Prop. 1 that enable computational benefits.
@@
$T_{t}(m,\alpha)=\mathbf{1}\left[\max_{1\leq i\leq t}K_{i}(m)\geq 1/\alpha\right].$ (6)
@@
For all $c\geq 1/4$, $T_{t}(m,\alpha)$ defines a sequential test with type-I error $\alpha$ for null hypothesis $m$, i.e., if $\mu=m$ (equivalently, $P_{1}\in\mathcal{P}(m)$), $\mathbb{P}(\exists t\in\mathbb{N}:T_{t}(m)=1)\leq\alpha$.
@@
The convexity of our confidence sequence $C_{t}(\alpha)$ follows directly from Proposition 1 (proof in Appendix A.5): at time $t=0$, $C_{0}(\alpha)=[0,1]$; for $t\in\mathbb{N}$, each test results in a convex set; repeated intersections $C_{t}(\alpha)$ is also convex. Note that this property does not generally hold in testing-by-betting schemes by Waudby-Smith and Ramdas, (2023).
Remark 1.
For example, the AGRAPA betting scheme, despite its similar form to our choice of betting scheme, is not guaranteed to form convex confidence sets (Waudby-Smith and Ramdas,, 2023, Appendix E.4). This requires one to test each $m\in[0,1]$ to construct confidence sets (or each $m\in R$ to determine whether one can reject a region of interest $R$). Our choice of $\lambda$ remedies this issue by (1) forming strictly convex confidence sets with respect to $m$, avoiding the need to test a grid of values over $[0,1]$ when constructing confidence sets, and thus (2) enabling a simple characterization of the minima $m$ in a convex region $R$, avoiding the need to test a fine grid of $m\in R$. These two benefits reduce the computational burden of our proposed testing method PEAK, a generalization of this simple betting scheme $K_{t}(m)$, discussed in Section 4.
While Theorems 1 and 2 guarantee that $\mu$ will be identified in our test, they do not provide additional insights. To better understand the behavior of our test, in Section 3.1, we characterize the asymptotic $e$-power/growth rate (Definition 2) of our process, $K_{t}(m)$, which defines the expected linear rate of increase of $l_{t}(m)$ as $t\rightarrow\infty$. We discuss the selection of parameter $c$ in Lemma 1. In Lemma 2, we establish that the worst-case instance of $e$-power is strictly positive for $m\neq\mu$. Finally, given that our test is nonparametric, we compare the asymptotic growth rate of our test in the parametric setting to the best-achievable growth rate in Lemma 3, establishing that the $e$-power of our test is near-optimal in the parametric setting.